OpenStudy (el_arrow):

for the given point in rectangular coordinates, find two sets of polar coordinates for the point (2sqrt(3),2) , 0

OpenStudy (el_arrow):

i got r = 4 how do i get theta?

OpenStudy (el_arrow):

@SolomonZelman

OpenStudy (jdoe0001):

right the coordinates are (x,y) to get the angle, recall that tangent = y/x

OpenStudy (el_arrow):

i know theta=tan^-1(2/2sqrt(3)) but i cant figure out the angles

OpenStudy (el_arrow):

is there an easy way to this?

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

nope... you'd need to get the arcTangent to get \(\theta\)

OpenStudy (jdoe0001):

hmmm oddly enough, is a known angle

OpenStudy (el_arrow):

would the angle be pi/3?

OpenStudy (jdoe0001):

so...that'd be \(\bf \cfrac{\pi }{6}\) now, keep in mind that arcTangent has a domain of \(\bf -\cfrac{\pi }{2}<\theta<\cfrac{\pi }{2}\)

OpenStudy (jdoe0001):

so the other angle, will be in the 4th quadrant

OpenStudy (el_arrow):

in the 4th quadrant.....would it be 5pi/6?

OpenStudy (jdoe0001):

well... 5/6 of \(\pi\) is not even one whole \(\pi\) so..to land on the 4th quadrant it has to be much more than that

OpenStudy (el_arrow):

i mean 11pi/6

OpenStudy (jdoe0001):

keep in mind that \(\bf \cfrac{6\pi }{6}\implies \pi \)

OpenStudy (jdoe0001):

yeap

OpenStudy (el_arrow):

okay so the two sets would be (4,pi/6) and (4,11pi/6)?

OpenStudy (jdoe0001):

yeap

OpenStudy (el_arrow):

alright thanks so much for your help @jdoe0001

OpenStudy (jdoe0001):

yw